OpenStudy (tw101):

Not sure where to start with this question : A sports store sells 70 pairs of shoes per day for $40 each. The owner estimates that for each $3 increase in price, 3 fewer sales are made each day. What price should be charged to maximize profit?

OpenStudy (tw101):

@phi

OpenStudy (phi):

try to write down equations

OpenStudy (tw101):

How? like 70 pairs a day = 70x? and then the cost of 40? like 70x=40? This is my first assignment on this so im just trying to figure it out

OpenStudy (phi):

That is the hard part... I would focus on $3 increase in price, 3 fewer sales which is equivalent to $1 up in price causes 1 fewer sale if we call the price increase x then we can say total = (40+x)(70-x) for x=0, we get the original conditions. we have to find the "x" value so that total is as big as possible

OpenStudy (tw101):

Okay, with you so far

OpenStudy (tw101):

so solve for x?

OpenStudy (phi):

once you have an equation, it's math. there are two ways: use calculus to find a min/max or in this case, the equation is a parabola, with a \( \cap \) shape, so the max value is its vertex

OpenStudy (tw101):

Okay ill put the equation into my caculator

OpenStudy (tw101):

I think I did something wrong

OpenStudy (tw101):

i put it in and got x=15 +- the square root of 3025-t @phi

OpenStudy (tw101):

or 55-t simplified

OpenStudy (phi):

what is t ?

OpenStudy (tw101):

the total

OpenStudy (tw101):

would the vertex be 15?

OpenStudy (phi):

if you use y instead of t (so it looks "normal") y= -x^2 +30x +2800

OpenStudy (tw101):

for x

OpenStudy (tw101):

oh okay

OpenStudy (tw101):

and the vertex for that equation would be -15 right?

OpenStudy (phi):

yes, the vertex is at -b/(2a) where a, b (and c) are the coefficients: y = a x^2 + b x + c matching that with y= -x^2 +30x +2800 a is -1 (-x^2 means -1 * x^2) b is 30 and -b/(2a) is -30/(2*-1) = 15

OpenStudy (phi):

so x=15 (not -15)

OpenStudy (tw101):

oh okay I see i second guessed myself

OpenStudy (tw101):

thank you so much for your help! its a little clearer now.

OpenStudy (phi):

you still have to answer the question: What price should be charged to maximize profit?

OpenStudy (tw101):

to find that out you'd substitute in 15 for x in the equation, right?

OpenStudy (phi):

that would give the total dollars. they want the cost of one pair

OpenStudy (phi):

if you look at the "start" equation total = (40+x)(70-x) which of those represents the cost of one pair ?

OpenStudy (tw101):

the 40+ x?

OpenStudy (phi):

yes

OpenStudy (tw101):

so the total would be 55?

OpenStudy (tw101):

no wait

OpenStudy (tw101):

sorry that doesnt make sense

OpenStudy (phi):

the price is 55 (and the number of pairs sold will be 70 - 15 = 55) and at the total profit would be 55*55 = 3025

OpenStudy (tw101):

oh really? okay haha i guess i should stop second guessing myself sorry

OpenStudy (tw101):

yes 3025 is what i got too

OpenStudy (tw101):

so the answer would be 55 then

OpenStudy (phi):

yes, What price should be charged to maximize profit? $55 per pair.

OpenStudy (tw101):

Okay, gotcha. Thank you so much!